# Deep Learning for Multivariate Time Series Imputation: A Survey

### v1

@@ Abstract
The ubiquitous missing values cause the multivariate time series data to be partially observed, destroying the integrity of time series and hindering the effective time series data analysis. Recently deep learning imputation methods have demonstrated remarkable success in elevating the quality of corrupted time series data, subsequently enhancing performance in downstream tasks. In this paper, we conduct a comprehensive survey on the recently proposed deep learning imputation methods. First, we propose a taxonomy for the reviewed methods, and then provide a structured review of these methods by highlighting their strengths and limitations. We also conduct empirical experiments to study different methods and compare their enhancement for downstream tasks. Finally, the open issues for future research on multivariate time series imputation are pointed out. All code and configurations of this work, including a regularly maintained multivariate time series imputation paper list, can be found in the GitHub repository https://github.com/WenjieDu/Awesome_Imputation.
@@ 1 Introduction
[…]
@@
More recently, deep learning imputation methods have shown great modeling ability in missing data imputation. These methods exploit powerful deep learning models like Transformers, Variational AutoEncoders (VAEs), Generative Adversarial Networks (GANs), and diffusion models to capture the intrinsic properties and potentially complex dynamics of time series. In this way, deep learning imputation methods can learn the true underlying data distribution from the observed data, so as to predict more reliable and reasonable values for the missing components. We note that there are several related surveys Khayati et al. (2020); Fang and Wang (2020) that primarily focus on statistical and machine learning imputation methods, but they offer only limited consideration of deep learning imputation methods. Considering that multivariate time series imputation is a crucial data preprocessing step for subsequent time series analysis, a thorough and systematic survey on deep multivariate time series imputation methods would significantly contribute to the advancement of the time series community.
@@
In this paper, we endeavor to bridge the existing knowledge gap by providing a comprehensive summary of the latest developments in deep learning methods for multivariate time series imputation (MTSI). First, we present a succinct introduction to the topic, followed by the proposal of a novel taxonomy, categorizing approaches based on two perspectives: imputation uncertainty and neural network architecture. Imputation uncertainty reflects confidence in imputed values for missing data, and capturing this involves stochastically generating samples and conducting imputations based on these varied samples Little and Rubin (2019). Accordingly, we categorize imputation methods into predictive ones, offering fixed estimates, and generative ones, which provide a distribution of possible values to account for imputation uncertainty. For neural network architecture, we explore a range of deep learning models tailored for MTSI, including Recurrent Neural Network (RNN)-based ones, Graph Neural Network (GNN)-based ones, Convolutional Neural Network (CNN)-based ones, attention-based ones, Variational AutoEncoder (VAE)-based ones, Generative Adversarial Network (GAN)-based ones, and diffusion-based ones. To provide practical imputation guidelines in real scenarios, we conduct extensive empirical studies that examine multiple aspects of deep multivariate time series imputation models, including imputation performance and improvement on downstream tasks like classification. To the best of our knowledge, this is the first comprehensive and systematic review of deep learning algorithms in the realm of MSTI, aiming to stimulate further research in this field. A corresponding resource that has been continuously updated can be found in our GitHub repository [^1] https://github.com/WenjieDu/Awesome_Imputation.
In summary, the contributions of this paper include: 1) A new taxonomy for deep multivariate time series imputation methods, considering imputation uncertainty and neural network architecture, with a comprehensive methodological review; 2) A thorough empirical evaluation of imputation algorithms via the PyPOTS toolkit we developed; 3) An exploration of future research opportunities for MTSI.
@@ 2.1 Background of MTSI
A complete time series dataset on $[0,T]$ typically can be denoted as $\mathcal{D}=\{\mathbf{X}_{i},\mathbf{t}_{i}\}_{i=1}^{N}$. Hereby, $\mathbf{X}_{i}=\{x_{1:K,1:L}\}\in\mathcal{R}^{K\times L}$ and $\mathbf{t}_{i}=({t_{1},\cdots,t_{L}})\in[0,T]^{L}$, where $K$ is the number of features and $L$ is the length of time series. In the missing data context, each complete time series can be split into an observed and a missing part, i.e., $\mathbf{X}_{i}=\{\mathbf{X}^{o}_{i},\mathbf{X}^{m}_{i}\}$. For encoding the missingness, we also denote an observation matrix as $\mathbf{M}_{i}=\{m_{1:K,1:L}\}$, where $m_{k,l}=0$ if $x_{k,l}$ is missing at timestamp $t_{l}$, otherwise $m_{k,l}=1$. Furthermore, we can also calculate a time-lag matrix $\boldsymbol{\delta}_{i}=\{\delta_{1:K,1:L}\}$ by the following rule,
@@
The missing mechanisms, i.e., the cause of missing data, represent the statistical relationship between observations and the probability of missing data Nakagawa (2015). In real-life scenarios, missing mechanisms are inherently complex, and the performance of an imputation model is significantly influenced by how closely the assumptions we make align with the actual missing data mechanisms. According to Robin’s theory Rubin (1976), the missing mechanisms fall into three categories: Missing Completely At Random (MCAR), Missing At Random (MAR), and Missing Not At Random (MNAR). MCAR implies that the probability of data being missing is independent of both the observed and missing data. Conversely, MAR indicates that the missing mechanism depends solely on the observed data. MNAR suggests that the missingness is related to the missing data itself and may also be influenced by the observed data. These three mechanisms can be formally defined as follows:
@@
2.2 Taxonomy of Imputation Methods
Figure 1: The taxonomy of deep learning methods for multivariate time series imputation from the view of imputation uncertainty and neural network architecture. For each category, one representative model is picked to display.
@@ 2.2 Taxonomy of Imputation Methods
To summarize the existing deep multivariate time series imputation methods, we propose a taxonomy from the perspectives of imputation uncertainty and neural network architecture as illustrated in Figure 1, and provide a more detailed summary of these methods in Table 1. For imputation uncertainty, we categorize imputation methods into predictive and generative types, based on their ability to yield varied imputations that reflect the inherent uncertainty in the imputation process. In the context of the neural network architecture, we examine prominent deep learning models specifically designed for time series imputation. The discussed models encompass RNN-based ones, CNN-based ones, GNN-based ones, attention-based ones, VAE-based ones, GAN-based ones, and diffusion-based ones. In the following two sections, we will delve into and discuss the existing deep time series imputation methods from these two perspectives.
@@ 3.4 GNN-based Models
GNN-based models, treating time series as graph sequences, reconstruct missing values using learned node representations. The authors in Cini et al. (2022) introduce GRIN, the first graph-based recurrent architecture for MTSI. GRIN utilizes a bidirectional graph recurrent neural network to effectively harness both temporal dynamics and spatial similarities, thereby achieving significant improvements in imputation accuracy. Furthermore, SPIN Marisca et al. (2022) is developed, integrating a unique sparse spatiotemporal attention mechanism into the GNN framework. This mechanism notably overcomes the error propagation issue of GRIN and bolsters robustness against the data sparsity presented by highly missing data.
@@ 3.5 Attention-based Models
Since Transformer is proposed in Vaswani et al. (2017), the self-attention mechanism has been widely used to model sequence data including time series Wen et al. (2023). CDSA Ma et al. (2019) is proposed to impute geo-tagged spatiotemporal data by learning from time, location, and measurement jointly. DeepMVI Bansal et al. (2021) integrates transformers with convolutional techniques, tailoring key-query designs to effectively address missing value imputation. For each time series, DeepMVI harnesses attention mechanisms to concurrently distill long-term seasonal, granular local, and cross-dimensional embeddings, which are concatenated to predict the final output. NRTSI Shan et al. (2023) directly leverages a Transformer encoder for modeling and takes time series data as a set of timestamp and measurement tuples. As a permutation model, this model has to iterate over the time dimension to process time series. SAITS Du et al. (2023) employs a self-supervised training scheme to deal with missing data, which integrates dual joint learning tasks: a masked imputation task and an observed reconstruction task. This method, featuring two diagonal-masked self-attention blocks and a weighted-combination block, leverages attention weights and missingness indicators to enhance imputation precision. In addition to the above models, the attention mechanism is also widely adapted to build the denoising network in diffusion models like CSDI Tashiro et al. (2021), MIDM Wang et al. (2023), PriSTI Liu et al. (2023), etc.
@@ 3.6 Pros and Cons
This subsection synthesizes the strengths and challenges of the predictive imputation methods discussed. RNN-based models, while adept at capturing sequential information, are inherently limited by their sequential processing nature and memory constraints, which may lead to scalability issues with long sequences Khayati et al. (2020). Although CNNs have decades of development and are useful feature extractors to capture neighborhood information and local connectivity, their kernel size and working mechanism intrinsically limit their performance on time series data as the backbone. Due to the attention mechanism, attention-based models generally outperform RNN-based and CNN-based methods in imputation tasks due to their superior ability to handle long-range dependencies and parallel processing capabilities. GNN-based methods provide a deeper understanding of spatio-temporal dynamics, yet they often come with increased computational complexity, posing challenges for large-scale or high-dimensional data.
@@ 4.2 VAE-based Models
VAEs employ an encoder-decoder structure to approximate the true data distribution by maximizing the Evidence Lower Bound (ELBO) on the marginal likelihood. This ELBO enforces a Gaussian-distributed latent space from which the decoder reconstructs diverse data points
@@ 4.3 GAN-based Models
In Luo et al. (2018), authors propose a two-stage GAN imputation method (GRUI-GAN), which is the first GAN-based method for imputing time series data. GRUI-GAN first learns the distribution of the observed multivariate time series data by a standard adversarial training manner, and then optimizes the input noise of the generator to further maximize the similarity of the generated and observed multivariate time series data. However, the second stage in GRUI-GAN needs a lot of time to find the best matched input vector, and this vector is not always the best especially when the initial value of the “noise” is not properly set. Then, an end-to-end GAN imputation model $E^{2}$ Luo et al. (2019) is further proposed, where the generator takes a denoising autoencoder module to avoid the “noise” optimization stage in GRUI-GAN. Meanwhile, authors in  Liu et al. (2019) propose a non-autoregressive multi-resolution GAN model (NAOMI), where the generator is assembled by a forward-backward encoder and a multiresolution decoder. The imputed data are recursively generated by the multiresolution decoder in a non-autoregressive manner, which mitigates error accumulation in scenarios involving high-missing and long sequence time series data. On the other hand, in Miao et al. (2021), authors propose USGAN, which generates high-quality imputed data by integrating a discriminator with a temporal reminder matrix. This matrix introduces added complexity to the training of the discriminator and subsequently leads to improvements in the generator’s performance. Furthermore, they extend USGAN to a semi-supervised model SSGAN, by introducing an extra classifier. In this way, SSGAN takes advantage of the label information, so that the generator can estimate the missing values, conditioned on observed components and data labels at the same time.
@@ 4.4 Diffusion-based Models
As an emerging and potent category of generative models, diffusion models are adept at capturing complex data distributions by progressively adding and then reversing noise through a Markov chain of diffusion steps. Distinct from VAE, these models utilize a fixed training procedure and operate with high-dimensional latent variables that retain the dimensionality of the input data.
CSDI, introduced in Tashiro et al. (2021), stands out as the pioneering diffusion model specifically designed for MTSI. Different from conventional diffusion models, CSDI adopts a conditioned training approach, where a subset of observed data is utilized as conditional information to facilitate the generation of the remaining segment of observed data. However, the denoising network in CSDI relies on two transformers, exhibiting quadratic complexity concerning the number of variables and the time series length. This design limitation raises concerns about memory constraints, particularly when modeling extensive multivariate time series. In response to this challenge, a subsequent work by Alcaraz and Strodthoff (2023) introduces SSSD, which addresses the quadratic complexity issue by replacing transformers with structured state space models Gu et al. (2022). This modification proves advantageous, especially when handling lengthy multivariate time series, as it mitigates the risk of memory overflow. Another approach CSBI, introduced in  Chen et al. (2023), improves the efficiency by modeling the diffusion process as a Schrodinger bridge problem, which could be transformed into computation-friendly stochastic differential equations.
Moreover, the efficacy of diffusion models is notably influenced by the construction and utilization of conditional information. MIDM Wang et al. (2023) proposes to sample noise from a distribution conditional on observed data’s representations in the denoising process, In this way, it can explicitly preserve the intrinsic correlations between observed and missing data. PriSTI Liu et al. (2023) introduces the spatiotemporal dependencies as conditional information, i.e., provides the denoising network with spatiotemporal attention weights calculated by the conditional feature for spatiotemporal imputation. Additionally, DA-TASWDM Xu et al. (2023) suggests incorporating dynamic temporal relationships, i.e. the varying sampling densities, into the denoising network for medical time series imputation.

### v2

@@ Abstract
Missing values are ubiquitous in multivariate time series (MTS) data, posing significant challenges for accurate analysis and downstream applications. In recent years, deep learning-based methods have successfully handled missing data by leveraging complex temporal dependencies and learned data distributions. In this survey, we provide a comprehensive summary of deep learning approaches for multivariate time series imputation (MTSI) tasks. We propose a novel taxonomy that categorizes existing methods based on two key perspectives: imputation uncertainty and neural network architecture. Furthermore, we summarize existing MTSI toolkits with a particular emphasis on the PyPOTS Ecosystem, which provides an integrated and standardized foundation for MTSI research. Finally, we discuss key challenges and future research directions, which give insight for further MTSI research. This survey aims to serve as a valuable resource for researchers and practitioners in the field of time series analysis and missing data imputation tasks.
@@ 1 Introduction
Figure 1: The framework of our survey.
The data collection process of multivariate time series in various fields is often fraught with difficulties and uncertainty. In IoT systems, sensor failures and unstable environments lead to missing measurements Li et al. (2023c). Clinical studies face challenges from irregular sampling and privacy concerns Ibrahim et al. (2012); Esteban et al. (2017); Qian et al. (2024). Financial and transportation systems encounter data gaps due to system downtime and communication issues Bai and Ng (2008); Gong et al. (2021). These missing values can significantly affect the accuracy and reliability of downstream analysis and decision-making. In real-world datasets like PhysioNet2012 Silva et al. (2012b), missing rates can exceed 80%. Consequently, exploring how to reasonably and effectively impute missing components in multivariate time series is attractive and essential.
@@
More recently, deep learning-based imputation methods have demonstrated strong modelling capabilities for handling missing data. These approaches leverage advanced architectures such as Transformers, Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), diffusion models, Pre-trained Foundation Models (PFMs), and Large Language Models (LLMs) to capture the underlying structures and complex temporal dynamics of time series data. By learning the true data distribution from observed values, deep learning imputation methods can generate more reliable and contextually appropriate estimates for missing components. While several surveys exist on imputation techniques Khayati et al. (2020); Fang and Wang (2020), they primarily focus on statistical and traditional machine learning approaches, offering limited discussion on deep learning-based methods. Given that multivariate time series imputation is a critical data preprocessing step for downstream time series analysis, a comprehensive and systematic survey on deep learning-driven imputation methods would provide valuable insights and contribute significantly to the research community.
@@
In this survey paper, we endeavor to bridge the existing knowledge gap by providing a comprehensive summary of the latest developments in deep learning methods for multivariate time series imputation (MTSI). The framework of this survey is illustrated in Figure 1. In detail, we first present a succinct introduction to the topic, followed by the proposal of a novel taxonomy, categorizing approaches based on two perspectives: imputation uncertainty and neural network architecture. Imputation uncertainty quantifies the confidence in estimated values for missing data. To capture this, multiple stochastic samples are generated, and imputations are performed across these variations Little and Rubin (2019). Accordingly, we categorize imputation methods into predictive ones, offering fixed estimates, and generative ones, which provide a distribution of possible values to account for imputation uncertainty. For neural network architecture, we explore a range of deep learning models tailored for MTSI, including RNN-based ones, GNN-based ones, CNN-based ones, attention-based ones, VAE-based ones, GAN-based ones, diffusion-based ones, PFM-based ones, and LLM-based ones.
To the best of our knowledge, this is the first comprehensive and systematic review of deep learning algorithms in the realm of MTSI, aiming to stimulate further research in this field. A corresponding resource that has been continuously updated can be found on our GitHub repository [^1] https://github.com/WenjieDu/Awesome_Imputation. In summary, the contributions of this survey include:
  1. 1.
    We introduce a novel taxonomy for deep multivariate time series imputation, and categorize them based on imputation uncertainty and neural network architecture.
  2. 2.
    We provide a comprehensive overview of existing MTSI toolkits. In particular, we highlight the PyPOTS Ecosystem, which integrates diverse imputation algorithms, standardized pipelines, and benchmarking resources, facilitating accessible and reproducible MTSI research.
  3. 3.
    We identify future directions, including missingness patterns, downstream task integration, and model scalability, offering insights to guide further advancements.
@@ 2.1 Background of MTSI
A complete time-series dataset on $[0,T]$ typically can be denoted as $\mathcal{D}=\{\mathbf{X}_{i},\mathbf{t}_{i}\}_{i=1}^{N}$. Hereby, $\mathbf{X}_{i}=\{x_{1:K,1:L}\}\in\mathcal{R}^{K\times L}$ and $\mathbf{t}_{i}=({t_{1},\cdots,t_{L}})\in[0,T]^{L}$, where $K$ is the number of features and $L$ is the length of time series. In the missing data context, each complete time series can be split into an observed and a missing part, i.e., $\mathbf{X}_{i}=\{\mathbf{X}^{o}_{i},\mathbf{X}^{m}_{i}\}$. For encoding the missingness, we also denote an observation matrix as $\mathbf{M}_{i}=\{m_{1:K,1:L}\}$, where $m_{k,l}=0$ if $x_{k,l}$ is missing at timestamp $t_{l}$, otherwise $m_{k,l}=1$. Furthermore, we can also calculate a time-lag matrix $\boldsymbol{\delta}_{i}=\{\delta_{1:K,1:L}\}$ by the following rule:
@@
The missing mechanisms, i.e., the cause of missing data, define the statistical relationship between observations and the probability of missingness Nakagawa (2015). In real-world scenarios, missing mechanisms are inherently complex, and imputation model performance heavily depends on how well our assumptions align with the actual missing data patterns. According to Rubin’s theory Rubin (1976), missing mechanisms fall into three categories: Missing Completely At Random (MCAR), Missing At Random (MAR), and Missing Not At Random (MNAR). MCAR implies that missingness is independent of both observed and missing data. Conversely, MAR indicates that missingness depends solely on observed data. MNAR suggests that missingness is related to the missing data itself and may also be influenced by observed data. These mechanisms can be formally defined as follows:
@@
2.2 Taxonomy of Deep Learning-based MTSI
Figure 2: The taxonomy of deep learning methods for multivariate time series imputation from the view of imputation uncertainty and neural network architecture.
@@ 2.2 Taxonomy of Deep Learning-based MTSI
[…]
To summarize the existing deep multivariate time series imputation methods, we propose a taxonomy from the perspectives of imputation uncertainty and neural network architecture as illustrated in Figure 2, and provide a more detailed summary of these methods in Table 1. Regarding large model-based methods, we separate them as a category considering their application strategies are quite different from general neural network algorithms. For imputation uncertainty, we categorize imputation methods into predictive and generative types, based on their ability to yield varied imputations that reflect the inherent uncertainty in the imputation process. In the context of the neural network architecture, we examine prominent deep learning models specifically designed for multivariate time series imputation. The discussed models encompass RNN-based ones, CNN-based ones, GNN-based ones, attention-based ones, VAE-based ones, GAN-based ones, and diffusion-based ones. In the following sections, we will delve into and discuss the existing deep time series imputation methods from these two perspectives.
@@ 3.4 GNN-based Models
GNN-based models treat time series as graph sequences, reconstructing missing values via learned node representations. Cini et al. (2022) introduces GRIN, the first graph-based recurrent architecture for MTSI, leveraging a bidirectional graph recurrent neural network to capture temporal dynamics and spatial similarities, significantly improving imputation accuracy. SPIN Marisca et al. (2022) further integrates a sparse spatiotemporal attention mechanism into the GNN framework, mitigating GRIN’s error propagation and enhancing robustness against data sparsity.
@@ 3.5 Attention-based Models
Since Transformer is proposed in Vaswani et al. (2017), the self-attention mechanism has been widely used to model sequence data including time series Wen et al. (2023). DeepMVI Bansal et al. (2021) integrates transformers with convolutional techniques, tailoring key-query designs to effectively address missing value imputation. For each time series, DeepMVI harnesses attention mechanisms to concurrently distill long-term seasonal, granular local, and cross-dimensional embeddings, which are concatenated to predict the final output. SAITS Du et al. (2023) employs a self-supervised training scheme to deal with missing data, which integrates dual joint learning tasks: a masked imputation task and an observed reconstruction task. This method, featuring two diagonal-masked self-attention blocks and a weighted-combination block, leverages attention weights and missingness indicators to enhance imputation precision. Besides, ImputeFormer Nie et al. (2024) introduces a novel Transformer-based framework that leverages self-attention and temporal context modeling to accurately recover missing values. In addition to the above models, the attention mechanism is also widely adapted to build the denoising network in diffusion models like CSDI Tashiro et al. (2021), MIDM Wang et al. (2023), PriSTI Liu et al. (2023a), Diffusion-TS Yuan and Qiao (2024) and in GNN-based models like Casper Jing et al. (2024) and HSPGNN Liang et al. (2024).
@@ 3.6 Pros and Cons
This subsection synthesizes the strengths and challenges of the predictive imputation methods discussed. RNN-based models, while adept at capturing sequential information, are inherently limited by their sequential processing nature and memory constraints, which may lead to scalability issues with long sequences Khayati et al. (2020). Although CNNs have decades of development and are useful feature extractors to capture neighborhood information and local connectivity, their kernel size and working mechanism intrinsically limit their performance on time-series data as the backbone. Due to the attention mechanism, attention-based models generally outperform RNN-based and CNN-based methods in imputation tasks due to their superior ability to handle long-range dependencies and parallel processing capabilities. GNN-based methods provide a deeper understanding of spatio-temporal dynamics, yet they often come with increased computational complexity, posing challenges for large-scale or high-dimensional data.
@@ 4.2 VAE-based Models
VAEs employ an encoder-decoder structure to approximate the true data distribution by maximizing the Evidence Lower Bound (ELBO) on the marginal likelihood. This ELBO enforces a Gaussian-distributed latent space from which the decoder reconstructs diverse data points.
@@ 4.3 GAN-based Models
In Luo et al. (2018), authors propose a two-stage GAN imputation method (GRUI-GAN), which is the first GAN-based method for imputing time-series data. GRUI-GAN first learns the distribution of the observed multivariate time-series data by a standard adversarial training manner and then optimizes the input noise of the generator to further maximize the similarity of the generated and observed multivariate time series. However, the second stage in GRUI-GAN needs a lot of time to find the best matched input vector, and this vector is not always the best especially when the initial value of the “noise” is not properly set. Then, an end-to-end GAN imputation model $E^{2}$ Luo et al. (2019) is further proposed, where the generator takes a denoising autoencoder module to avoid the “noise” optimization stage in GRUI-GAN. Meanwhile, authors in  Liu et al. (2019) propose a non-autoregressive multi-resolution GAN model (NAOMI), where the generator is assembled by a forward-backward encoder and a multiresolution decoder. The imputed data are recursively generated by the multiresolution decoder in a non-autoregressive manner, which mitigates error accumulation in scenarios involving high-missing and long sequence time series. On the other hand, in Miao et al. (2021), authors propose USGAN, which generates high-quality imputed data by integrating a discriminator with a temporal reminder matrix. This matrix introduces added complexity to the training of the discriminator and subsequently leads to improvements in the generator’s performance. Furthermore, they extend USGAN to a semi-supervised model SSGAN, by introducing an extra classifier. In this way, SSGAN leverages label information, allowing the generator to estimate missing values while conditioning on observed components and labels simultaneously.
@@ 4.4 Diffusion-based Models
As an emerging and potent category of generative models, diffusion models are adept at capturing complex data distributions by progressively adding and then reversing noise through a Markov chain of diffusion steps. Distinct from VAE, these models utilize a fixed training procedure and operate with high-dimensional latent variables that retain the dimensionality of the input data Yang et al. (2024b).
CSDI, introduced in Tashiro et al. (2021), stands out as the pioneering diffusion model specifically designed for MTSI. Different from conventional diffusion models, CSDI adopts a conditioned training approach, where a subset of observed data is utilized as conditional information to facilitate the generation of the remaining segment of observed data. However, the denoising network in CSDI relies on two transformers, exhibiting quadratic complexity concerning the number of variables and the time series length. This design limitation raises concerns about memory constraints, particularly when modeling extensive multivariate time series. In response to this challenge, a subsequent work by Alcaraz and Strodthoff (2023) introduces SSSD, which addresses the quadratic complexity issue by replacing transformers with structured state space models Gu et al. (2022). This modification proves advantageous, especially when handling lengthy multivariate time series, as it mitigates the risk of memory overflow. Another approach CSBI, introduced in  Chen et al. (2023c), improves the efficiency by modeling the diffusion process as a Schrodinger bridge problem, which could be transformed into computation-friendly stochastic differential equations. Also, SADI Dai et al. (2024) is a similarity-aware diffusion model that leverages a self-attention mechanism to capture inter-patient similarities for effective imputation of missing values. MTSCI Zhou et al. (2024) leverages cross-channel correlations and multi-scale temporal dynamics features to effectively recover missing values.
Moreover, the efficacy of diffusion models is notably influenced by the construction and utilization of conditional information. MIDM Wang et al. (2023) proposes to sample noise from a distribution conditional on observed data’s representations in the denoising process, In this way, it can explicitly preserve the intrinsic correlations between observed and missing data. PriSTI Liu et al. (2023a) introduces the spatiotemporal dependencies as conditional information, i.e., provides the denoising network with spatiotemporal attention weights calculated by the conditional feature for spatiotemporal imputation. Besides, FGTI Yang et al. (2024a) is a frequency-guided framework that leverages spectral analysis to effectively capture both global periodic patterns and local temporal dynamics, thereby enhancing missing data recovery.